

********************************************************************
UARS WEEKLY STATUS REPORT - 30 OCT 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occulation
data during this past week without apparent problems.
Reprocessing of all SUSIM experiment days to version 20 continues.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE continues to operate successfully with all engineering parameters
within nominal operating range. HALOE science data results continues to be
excellent. HALOE will be turned off at the end of day today GMT in
preparation for the upcoming yaw.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI completed a fulltime-on observation period ending Oct. 28. We are now
in daytime-only mode. Operations are as expected.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues with no
problems. The MPG continues to assist the Project Scientist with the
scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
53 days HAL (4.80) PROD (L1) (1873R-1845C-24F=4L)
0 days HAL (4.80) PROD (L3) (1873R-1139C=734L)
63 days SOL (4.62) PROD (L2-L3) (2349R-1914C=435L)
65 days SUS (7.00) PROD (L1-L2) (300R-300C=0L)
149 days SUS (7.00) PROD (L1-L3) (1249R-328C=921L)
3 days CLA (7.00) ENG (L1-L3) (18R-3C-2F=13L)
CDHF2 is once again stable. We have resumed NRT, QL and Level 0 receipt
through CDHF2 which is the primary system on 10/26/98 rather than UARSMS
which is the backup.
The CDHF loaded August 1998 DAAC data.
CDHF pending software upgrades: IDL 5.1, OpenIngres, UCSS 10.1, MSS 4.1,
NRT 2.5, UOAS 5.4
***********
FOT
***********
REPORTING PERIOD
This report covers 17 October 1998 (Orbit 38803, GMT Day 290) through 23
October 1998 (Orbit 38893, GMT Day 296).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 290/0025 38789 On
WINDII 295/2317 38878 Off
The beta angle ranged between 47.6 degrees to 38.1 degrees for this report
period. The beta angle is now increasing toward a minimum angle of 3.0
degrees on 02 November 1998 (DOY 306).
Flight Operations continues to work with the PEM instrument to fine tune
the operational command sequences associated with the OBC Power Monitor
day-only operational capability.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 16 October 1998 and 23 October 1998 are provided in
Appendix D.
The clock error ranged from 11.4 to -11.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% PEM 02
% WINDII 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 09
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/210 - 2287.496941
Transponder B: 98/211 - 2287.499161
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/295 show the observatory data loss to be
21 hours, 10 minutes, 35.060 seconds (an increase of 1 minute and 2.72
seconds since last report period). This is a 0.0341 percent data loss
which equals a 99.9659 percent data capture for the mission.
The increase of 1 minute and 2.72 seconds occurred over six (6) days:
7 0.672 seconds on Day 289 due to hits in data;
7 3.360 seconds on Day 291 due to hits in data;
7 1.184 seconds on Day 292 due to hits in data;
7 1.888 seconds on Day 293 due to hits in data;
7 52.544 seconds on Day 294 due to IP anomaly (data hits) and
insufficient overlap; and
7 3.072 seconds on Day 295 due to hits in data.
The FOT modified versions of several CMS resident Relative Time Sequences
(RTSs) to facilitate proper PMON-controlled day-only operation of the PEM
instrument.
CMS performance problems continue but does not cause any operational
concerns.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
Investigation continues into possible impacts of the Leonid meteorite
shower in November 1998.
Investigating the feasibility of a new PMON action to change battery V/T
curves based on loading to mitigate battery risks in sun pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - No change.
Multi Satellite Operations Control Center (MSOCC) - No change.
Command Management System (CMS) - The FOT received a valid Y2K 3-day S/C
Orbital Events file from FDF this week. Upon receipt of this file, the FOT
attempted to complete the remainder of the seventeen (17) Y2K test
functions which have not been completed to date. The FOT was able to
successfully complete all of the test functions except the one related to
RSSPP SchedulingS. Apparently, the CMS requires the existence of the
SC_YAWMAN mnemonic in the S/C Orbital Events file to determine spacecraft
direction for SSPP Scheduling. Since the FOT only received a 3-day Orbital
Events file (and never received an 8-week Orbital Events file to use as a
long term baseline), there is no SC_YAWMAN mnemonic for the CMS to
reference. The missing SC_YAWMAN mnemonic results in a FATAL error during
SSPP Scheduling. The FOT and the test team will discuss possible solutions
to this problem early next week.
The only other test function that has not been tested to date is the one
related to Star Catalog processing. The FOT is still waiting for a Star
Catalog file from FDF which has an epoch time close enough (or into) year
2000 to allow valid CMS processing of that file.
Flight Dynamics Facility (FDF) - Received a set of 18-month planning aids
from FDF this week. Specifically, the FOT received a Brouwer Mean Orbital
Events file, a Target Visibility file, and an Orbit Parameter Profile file.
IP NETWORK TESTING
Space Network (SN) - Operational.
NCC 98 system - In planning.
Deep Space Network (DSN) and Ground Spaceflight Tracking and Data Network
(GSTDN) - No change. Attempting to coordinate test with NIA personnel.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
GSFC Propulsion Branch reported that the propulsion engineers will be ready
to support the UARS simulated line refill test after 04 November 1998. This
will be the transition from their previous assignment to a new assignment,
so there is a limited window of availability. FOT personnel are preparing
the commanding sequences to achieve a 16 ms and a 8 ms command timing
between ON and OFF of the latch valve. This will be checked with the
simulator before using the latch valve test hardware. Jackson and Tull
were funded to built the test hardware but do not have funds to support
initial checkout operation or operation of the test.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Flight operations is providing assistance on both sustaining engineering
and re-engineering activities to incorporate needed features into the
Trending system. Items discussed included pseudo-telemetry, statistical
operators (also known as "UGPL command line options"), and conditional data
filters.
MEETINGS
The next battery telecon with JPL scheduled for 20 October 1998 at 1:00 pm
was posponed until 27 October 1998.
The FOT attended a meeting on 23 October 1998 with representatives from the
Y2K CMS test team, FDF, and FORMATS development team to discuss outstanding
Y2K test issues. Topics of discussion were the generation and delivery of
a Y2K Star Catalog as well as the FOTUs necessity for a CMS version of the
Alignment and Calibration file. FDF is still attempting to generate a
valid Star Catalog, and the FOT has determined that a printout of the
Alignment and Calibration data (versus a CMS compatible file) is
acceptable.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 02
November 1998.
UARS will pass through the Leonid meteor shower on 17 November 1998 at
19:43 UTC (+/- 3 hours).
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
Erik Larson, UARS System Engineer, submitted his resignation effective 01
November 1998.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 23 October 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
38789 290/0025 98-401 WINDII Power ON
38820 292/0220 98-402 OBC Clk Adj fr 20.6 to 0.00 msec/day
38845 293/1825 98-403 OBC Clk Adj fr 0.00 to 20.6 msec/day
38858 294/1506 98-404 PEM Changed HEPS Turn-On Timing
38878 295/2317 98-405 WINDII Power OFF
38882 296/0550 98-406 OBC Clk Adj fr 20.6 to 0.00 msec/day
38889 296/1655 98-407 PMON/PEM PEM Turn-on test
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 23 October 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 23 October 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
38820 292/02:38 (19 Oct 98) from 20.6 to 00.0 msec/day
38845 293/18:31 (20 Oct 98) from 00.0 to 20.6 msec/day
38882 296/05:55 (23 Oct 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 23 October 1998
16 Oct 1998 - GMT Day 289 Beta = 47.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 17.9 0.000 0.0
2 +22.4/-11.2 +5.15/+3.87 +27.2/-14.8 26.1 1.028 20.4
3 +67.2/-11.2 +3.87/+2.60 +26.4/-14.4 26.1 1.026 19.9
23 Oct 1998 - GMT Day 296 Beta = 38.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-3.99 +0.0/ +0.0 17.9 0.000 0.0
2 +44.8/-5.6 +5.47/+4.51 +29.2/-14.4 26.1 1.028 20.6
3 +134.4/0.0 +4.51/+3.23 +28.4/-14.4 26.1 1.026 20.2
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 23 October 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-073 FOT 38820 292/0226 N/A TAC 8 Crashed
98-074 FOT 38829 292/1456 N/A TAC 4 Crashed
98-075 FOT 38829 292/1609 N/A TAC 8 Crashed
98-076 FOT 38834 293/0038 N/A TAC 4 Crashed
98-077 FOT 38839 293/0904 N/A TAC 8 Crashed
98-078 FOT 38863 294/2303 20287 IP Anomaly
98-079 FOT 38864 295/0056 20288 IP Anomaly
98-080 FOT 38875 295/1759 N/A TAC 2 Crashed
98-082 FOT 38888 296/1514 20288 Drop-out/IP Anomaly
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-073 FOT 38820 292/02:27 TAC 8 CRASH
98-074 FOT 38828 292/14:56 TAC 4 CRASH
09=075 FOT 38829 292/16:09 TAC 8 CRASH
98-076 FOT 38834 293/00:38 TAC 4 CRASH
97-077 FOT 38839 293/09:04 TAC 2 CRASH
98-078 FOT MANY 294/ 20287 IP ANOMALY
98-079 FOT LISTED 295/ 20288 IP ANOMALY
98-080 FOT 38875 295/17:59 TAC 2 CRASH
98-081 FOT 38824 292/09:00 GND ERR, LATE ACQ
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-066 FOT Oct 98 See List 14170 GENERIC LATE ACQ
SEVNTFW43